OpenStudy (anonymous):

Average tuition at two-year colleges was $2361 in 2007, up from $2275 in 2006. What percent increase in coast was this?

OpenStudy (anonymous):

Take $2361-$2275=$86 cost increase then take $86/2275=0.037802198*100=3.7802198% increase